# SpecEdge: Scalable Edge-Assisted Serving Framework for Interactive LLMs

### v1

@@ Abstract
Large language models (LLMs) power many modern applications, but serving them at scale remains costly and resource-intensive. Current server-centric systems overlook consumer-grade GPUs at the edge. We introduce SpecEdge, an edge-assisted inference framework that splits LLM workloads between edge and server GPUs using a speculative decoding scheme, exchanging only token outputs over the network. SpecEdge employs proactive edge drafting to overlap edge token creation with server verification and pipeline-aware scheduling that interleaves multiple user requests to increase server-side throughput. Experiments show SpecEdge enhances overall cost efficiency by 1.91x through achieving 2.22x server throughput, and reduces inter token latency by 11.24% compared to a server-only baseline, introducing a scalable, cost-effective paradigm for LLM serving. [^1] The source code will be publicly available after the review.
@@ 1 Introduction
Large language models (LLMs) have become integral to modern applications such as conversational AI, code generation, and real-time content creation (Dubey et al., 2024; Lozhkov et al., 2024; Achiam et al., 2023; Touvron et al., 2023; Jiang et al., 2023; Brown et al., 2020). However, scaling LLM deployments to meet growing demand remains challenging when balancing operational costs against latency requirements.
A compelling opportunity exists to dramatically reduce LLM serving costs by leveraging consumer-grade GPUs at the network edge. The GeForce RTX 4090 (Nvidia, 2025a) delivers up to 330.3 TFLOPS for FP16 tensor operations with FP16 accumulate (Nvidia, 2024a), exceeding the 312 TFLOPS of the data-center-class A100 (Nvidia, 2025b), at 14.43x lower cost (GCP, 2025; Vas, 2025). With the widespread availability of these powerful edge devices (Valve, 2024; Nvidia, 2024b), an edge-assisted inference approach that offloads computation to these cost-effective resources could fundamentally transform the economics of LLM deployment.
Despite this opportunity, existing inference architectures fail to leverage these edge resources effectively. Current parallelization techniques (Shoeybi et al., 2019; Rasley et al., 2020; Aminabadi et al., 2022) that split computation within data centers break down over public internet conditions, where high latency and limited bandwidth make frequent communication of intermediate results impractical. Mainstream approaches like tensor and pipeline parallelism rely on high-bandwidth, low-latency interconnects such as NVLink or InfiniBand (Nvidia, 2025c, d). In wide-area networks (WANs), transferring intermediate model states between edge and server GPUs quickly becomes prohibitive, preventing meaningful collaboration between these heterogeneous resources.
@@ 2 Background and Related Work
LLM serving systems. Modern LLM frameworks address latency, throughput, and resource efficiency challenges through various optimizations. DeepSpeed-Inference (Rasley et al., 2020) and TensorRT-LLM (Nvidia, 2024c) leverage low-level GPU optimizations, model parallelism, and quantization, while vLLM (Kwon et al., 2023) introduces PagedAttention for efficient memory management. Modern parallelization strategies (Shoeybi et al., 2019; Aminabadi et al., 2022) enhance the performance with multiple GPUs. However, these approaches depend on data-center GPUs connected via specialized interconnects (InfiniBand, NVLink) with throughput exceeding hundreds of GB/s (Nvidia, 2025c)—speeds unattainable over wide area networks (WANs).
Speculative decoding. Speculative decoding (Leviathan et al., 2023; Chen et al., 2023) reduces latency by having a smaller auxiliary model generate multiple candidate tokens for parallel verification by the main model. The process involves three phases: drafting candidates, verification, and reconciliation for generated tokens. Recent advances have focused on more efficient drafting methods, either using lighter auxiliary models or the target model with reduced parameters (Bhendawade et al., 2024; Cai et al., 2024; Li et al., 2024; Stewart et al., 2024; Zhang et al., 2023).
Tree-based speculative decoding. Standard speculative decoding suffers from exponentially declining acceptance rates as sequence length increases (Leviathan et al., 2023; Chen et al., 2023). Tree-based approaches (Miao et al., 2024; Chen et al., 2024; Svirschevski et al., 2024; Cai et al., 2024; Sun et al., 2024) address this by exploring multiple paths simultaneously. Sequoia (Chen et al., 2024) and SpecExec (Svirschevski et al., 2024) further optimize by pruning unpromising branches.
Distributed LLM serving. Split-inference approaches like Petals (Borzunov et al., 2024) and Helix (Mei et al., 2024) distribute model layers and pipeline requests across multiple devices, improving throughput over memory offloading methods (Ren et al., 2021; Pudipeddi et al., 2020), but introduce network delays that increase latency compared to data center solutions. Several works (Timor et al., 2025; Liu et al., 2025; McDanel, 2024) utilize multiple GPU devices within a server node to overlap draft and verification tasks of speculative decoding for faster inference, yet with an exchange of higher cost as they require additional server devices per query.
On-device LLM inference. Frameworks like MLC LLM (MLC team, 2023-2024) and Web LLM enable fully on-device inference with smaller or quantized models. While recent approaches (Svirschevski et al., 2024; Song et al., 2024; Xue et al., 2024) showcase the potential of leveraging user devices, they compromise output quality and latency. In contrast, our approach retains the high-quality verification stage on the server—achieving a hybrid solution that outperforms purely centralized or fully local inference.
@@ 3 Problem and Motivation
Conventional split computing. One approach to utilizing edge GPUs is split computation across edge and server resources. Split computing has been extensively studied in machine learning literature (Kang et al., 2017; Zhou et al., 2019; Wang et al., 2020), predating transformer architectures. Traditionally, this approach divides neural network layers between server and edge, reducing server computational load or minimizing network communication by transmitting intermediate tensors rather than raw input data. However, this method is poorly suited for LLM serving due to the unique characteristics of transformers (Vaswani, 2017):
Limitation #1: Excessive latency. While layer-wise splitting has been used in distributed peer-to-peer systems (Borzunov et al., 2024; Mei et al., 2024) to enable serving large models otherwise infeasible on consumer GPUs, these approaches prioritize feasibility over performance. This is because, unlike earlier ML applications (He et al., 2016; Redmon, 2016; Diwan et al., 2023) that typically involve only a few communication rounds, LLM applications require communication for every generated token due to their autoregressive nature. In our distributed network setting (§5.3), layer-wise splitting increases latency by 2.35x compared to the server-only solution.
Limitation #2: Ineffectiveness in reducing I/O. LLM inference is inherently constrained by memory I/O (Pope et al., 2023), as generating a single token requires accessing the entire model’s parameters. Transformer architectures, with operations like general matrix-vector multiplication (GEMV), have low arithmetic intensity relative to their memory I/O demands. While layer-wise split computing reduces computational load by offloading layers to the edge, it fails to alleviate the I/O-bound nature of inference. A more effective approach is batch-verifying multiple candidate tokens, which significantly reduces the I/O overhead per token and increases GPU throughput.
@@ 4.1 Disaggregated LLM Decoding
In our novel design, edge GPUs generate candidate tokens and send them to the server, which verifies them in a single forward pass. The server returns both the verified tokens and one additional token to the edge device, which then updates its sequence, KV cache, and continues drafting. This cycle repeats until an end-of-sequence token is generated. Critically, this approach preserves the exact output distribution of the server model (Leviathan et al., 2023)—even when using different models for drafting and verification, the final output is guaranteed to be sampled from the same distribution as if generated by the server model alone.
@@ 4.2 Proactive Draft Generation at the Edge
Initial drafting. SpecEdge employs state-of-the-art tree-based drafting  (Svirschevski et al., 2024). However, SpecEdge is designed to be future-proof and not tied to any specific drafting technique. It simply requires a draft phase that produces candidate tokens and is compatible with various speculative decoding approaches, including tree-based methods (Chen et al., 2024; Miao et al., 2024), lossy/lossless methods, and earlier single-sequence schemes (Leviathan et al., 2023; Chen et al., 2023). As the field evolves, any advances in speculative decoding can be seamlessly integrated.
@@ 5 Evaluation
Implementation and Setup. Our system’s edge-assisted configuration utilizes a server-side NVIDIA A100 GPU connected to multiple edge-side NVIDIA RTX4090 GPUs over a wide-area network. The number of RTX4090 GPUs scales with the number of concurrent requests (batch size x 2). In our experiments, we measured an average round-trip time (RTT) of 14.07ms between the local edge node and our Google Cloud instance. We conducted evaluations across various models and datasets under diverse operating conditions.
Baseline and metrics. Our primary baseline is a server-only configuration employing tree-based speculative decoding, supplemented by autoregressive decoding and a layer-split approach that offloads part of the LLM’s layers to an edge device. SpecEdge can leverage either client-side or edge GPUs; in this evaluation, we assume it uses consumer-grade GPUs from edge cloud providers. Based on provider pricing (GCP, 2025; Vas, 2025), the server-side A100 40GB GPU costs \$4.05 per hour, while running the 32B model requires an A100 80GB at \$5.05 per hour. In comparison, SpecEdge adds \$0.35 per hour for each RTX 4090 used. Our key metrics include cost efficiency (generated tokens per dollar), server throughput (generated tokens per unit time) and inter token latency (user-perceived output latency). All configurations produce identical output distributions as they use the same underlying models.
Models and data sets. We use four different LLMs: Qwen3-32B/14B (Team, 2025), Vicuna-33B (Chiang et al., 2023) and Llama2-13B-chat-hf (Touvron et al., 2023). Unless specifically noted, all models are configured with a temperature setting of 0.7. For the draft models, we use five different models: Qwen3-1.7B/0.6B (Team, 2025), Sheared Llama-1.3B (Xia et al., 2023), Tiny Llama-1.1B (Zhang et al., 2024), and JackFram-160M (Miao et al., 2024). Finally, we use SpecBench (Xia et al., 2024), C4 (en) (Raffel et al., 2020), OpenAssistant conversations datasets (Köpf et al., 2024), WikiText-2 (Merity et al., 2016), and MTBench (Zheng et al., 2023).
@@
Qwen3 14B/1.7B Multi-turn bench 3.92±$\pm 1.51$ 3.98±$\pm 1.57$ 31.78 66.54 (2.09x) 28.25 50.60 (1.79x)
Translation 3.95±$\pm 1.47$ 4.25±$\pm 1.45$ 32.24 65.25 (2.02x) 28.66 49.47 (1.73x)
Summarization 3.73±$\pm 1.60$ 3.95±$\pm 1.61$ 29.70 67.53 (2.27x) 26.40 51.22 (1.94x)
QA 3.42±$\pm 1.57$ 3.59±$\pm 1.56$ 27.30 62.04 (2.27x) 24.26 47.09 (1.94x)
Math. 4.10±$\pm 1.48$ 4.25±$\pm 1.49$ 32.84 72.93 (2.22x) 29.19 55.28 (1.89x)
RAG 3.73±$\pm 1.53$ 3.83±$\pm 1.56$ 29.89 64.04 (2.14x) 26.57 48.78 (1.84x)
Qwen3 14B/0.6B Multi-turn bench 3.87±$\pm 1.41$ 4.41±$\pm 2.25$ 33.45 69.58 (2.08x) 29.73 52.97 (1.78x)
Translation 3.79±$\pm 1.48$ 4.67±$\pm 2.34$ 32.88 69.00 (2.10x) 29.22 52.23 (1.79x)
Summarization 3.68±$\pm 1.49$ 4.21±$\pm 2.16$ 31.17 68.60 (2.20x) 27.71 52.16 (1.88x)
QA 3.33±$\pm 1.41$ 3.79±$\pm 1.94$ 28.89 61.90 (2.14x) 25.68 46.98 (1.83x)
Math 3.90±$\pm 1.57$ 5.27±$\pm 2.27$ 33.53 83.88 (2.50x) 29.80 63.56 (2.13x)
RAG 3.53±$\pm 1.52$ 4.29±$\pm 2.16$ 30.07 69.51 (2.31x) 26.73 52.76 (1.97x)
Qwen3 32B/1.7B Multi-turn bench 4.22±$\pm 1.93$ 4.71±$\pm 2.66$ 24.96 56.47 (2.26x) 17.80 35.38 (1.99x)
Translation 4.08±$\pm 1.97$ 5.24±$\pm 2.84$ 24.33 58.79 (2.42x) 17.34 36.83 (2.12x)
Summarization 4.19±$\pm 2.01$ 4.52±$\pm 2.68$ 24.33 54.07 (2.42x) 17.65 33.90 (2.12x)
QA 3.62±$\pm 1.95$ 3.93±$\pm 2.49$ 21.59 46.14 (2.14x) 15.39 28.99 (1.88x)
Math. 4.60±$\pm 1.93$ 5.40±$\pm 2.78$ 27.52 64.01 (2.33x) 19.62 40.18 (2.05x)
RAG 3.89±$\pm 2.05$ 4.19±$\pm 2.69$ 22.67 49.46 (2.18x) 16.16 31.05 (1.92x)
@@ 5.1 End-to-end Performance and Cost-efficiency
Beyond throughput improvements, Figure 6 demonstrates that SpecEdge also reduces inter-token latency by an average of 11.24% compared to server-only speculative decoding across various batch sizes. This dual benefit—higher throughput with lower latency—is particularly valuable for interactive LLM applications where both resource efficiency and user experience are critical.
We extend our evaluation to other models, using Vicuna-33B and Llama2-13B-Chat as target LLMs across diverse datasets including C4, OAsst, WikiText-2, and MTBench. The results show that proactive edge drafting SpecEdge consistently improves performance, with greater benefits observed when the drafting models generate deeper drafts with better alignment toward the target LLMs. This alignment quality directly correlates with improvements in server throughput and reduced inter-token latency. The complete results for all combinations of models and sets are provided in the Appendix D.
@@ 5.2 Component-wise Benefit
However, simple disaggregation alone can increase latency and leave the server underutilized (§4.1). Figure 9 compares three progressive implementations with a 14B/1.7B model pair: basic disaggregation, disaggregation with proactive edge drafting (§4.2), and complete SpecEdge with pipeline-aware scheduling (§4.3). Basic disaggregation achieves only 32.76 tokens/s throughput with higher latency, while adding proactive drafting reduces inter-token latency to 28 ms. The complete SpecEdge with pipeline-aware scheduling dramatically increases server throughput to  67.89 tok/s (2.07× improvement) while maintaining latency, demonstrating the complementary benefits of each component.
@@ 5.3 System Sensitivity Analysis
Network RTT sensitivity. We evaluate the average inter token latency of SpecEdge against layer-split and server-only configurations across varying network round-trip times (RTTs), using Qwen32B on SpecBench. Layer-split configuration runs autoregressive decoding, where one-quarter of model layers run on an edge RTX4090, with the remainder on a server-side A100. The server-only configuration, unaffected by network RTT, runs tree-based speculative decoding entirely on an A100. Figure 12 shows that SpecEdge provides lower inter-token latency than the server-only baseline below 50 ms RTT, with a 13.90% gain at 15 ms RTT (36.47 ms vs. 42.36 ms). Even at 65 ms, SpecEdge ’s latency rises by only 22.00% over its 15 ms RTT performance (to 44.47 ms), remaining competitive. By contrast, layer-split is much slower: at 15 ms RTT, it is 2.73× slower than SpecEdge (99.63 ms vs. 36.47 ms), increasing to 3.35× slower performance at 50 ms RTT (134.90 ms vs. 40.26 ms). This resilience stems from SpecEdge ’s less frequent communication rounds and proactive edge drafting, which offset network latency more effectively than layer-split approaches.
@@
Performance with varying edge devices. Figure 12 presents the speedup achieved by SpecEdge when the server is assisted by different edge devices (RTX 4090, 4070 Ti Super, 3090, and 2080Ti) across three target-draft model combinations. The speedup is measured relative to default autoregressive decoding using only the A100 server GPU. The consistent speedup across all model combinations confirms the architecture’s robustness to different hardware configurations. As expected, more powerful edge GPUs like the RTX 4090 deliver greater speedups, while even more affordable options like the RTX 2080Ti still provide significant acceleration. This demonstrates that SpecEdge’s approach remains effective across a spectrum of edge hardware capabilities, allowing deployment flexibility.
@@
Performance with alternative drafting approaches. We also evaluated SpecEdge with a non-tree speculative decoding approach (Leviathan et al., 2023) to demonstrate versatility beyond tree-structured methods. Using this alternative architecture, SpecEdge achieved up to 1.96× higher server throughput (Llama2 13B/TL 1.1B pairing) and 1.67× better cost efficiency compared to server-only deployments. Performance gains remained consistent across different model combinations, with even the smallest draft model (JF 160M) delivering a 1.52× throughput improvement while preserving end-user speedup. Complete results are available in Appendix E.
@@
Detailed Case Study. Complementing our quantitative evaluation, Appendix F offers a visualization of SpecEdge in action. Using Llama models responding to a query about Dyson Spheres, we trace the complete token generation lifecycle—from initial drafting through verification and subsequent accelerated generation. The case study specifically highlights two key operational advantages: (1) how edge GPUs remain productive during server verification phases through proactive expansion, and (2) how successful draft alignments lead to deeper draft candidates in subsequent rounds.

### v2

@@ Abstract
Large language models (LLMs) power many modern applications, but serving them at scale remains costly and resource-intensive. Current server-centric systems overlook consumer-grade GPUs at the edge. We introduce SpecEdge, an edge-assisted inference framework that splits LLM workloads between edge and server GPUs using a speculative decoding scheme, exchanging only token outputs over the network. SpecEdge employs proactive edge drafting to overlap edge token creation with server verification and pipeline-aware scheduling that interleaves multiple user requests to increase server-side throughput. Experiments show SpecEdge enhances overall cost efficiency by 1.91× through achieving 2.22× server throughput, and reduces inter token latency by 11.24% compared to a server-only baseline, introducing a scalable, cost-effective paradigm for LLM serving. The code is available at https://github.com/kaist-ina/specedge
@@ 1 Introduction
Large language models (LLMs) have become integral to modern applications such as conversational AI, code generation, and real-time content creation (dubey2024llama; lozhkov2024starcoder; achiam2023gpt; touvron2023llama; jiang2023mistral; brown2020language). However, scaling LLM deployments to meet growing demand remains challenging when balancing operational costs against latency requirements.
A compelling opportunity exists to dramatically reduce LLM serving costs by leveraging consumer-grade GPUs at the network edge. The GeForce RTX 4090 (RTX4090_spec) delivers up to 330.3 TFLOPS for FP16 tensor operations with FP16 accumulate (Nvidia_ada_GPU), exceeding the 312 TFLOPS of the data-center-class A100 (A100_spec), at 14.43x lower cost (GCP; Vastai). With the widespread availability of these powerful edge devices (steam_hardware_survey; Nvidia_Financial_Reports), an edge-assisted inference approach that offloads computation to these cost-effective resources could fundamentally transform the economics of LLM deployment.
Despite this opportunity, existing inference architectures fail to leverage these edge resources effectively. Current parallelization techniques (shoeybi2019megatron; rasley2020deepspeed; aminabadi2022deepspeed) that split computation within data centers break down over public internet conditions, where high latency and limited bandwidth make frequent communication of intermediate results impractical. Mainstream approaches like tensor and pipeline parallelism rely on high-bandwidth, low-latency interconnects such as NVLink or InfiniBand (NVLink; InfiniBand). In wide-area networks (WANs), transferring intermediate model states between edge and server GPUs quickly becomes prohibitive, preventing meaningful collaboration between these heterogeneous resources.
@@ 2 Background and Related Work
LLM serving systems. Modern LLM frameworks address latency, throughput, and resource efficiency challenges through various optimizations. DeepSpeed-Inference (rasley2020deepspeed) and TensorRT-LLM (TensorRT-LLM) leverage low-level GPU optimizations, model parallelism, and quantization, while vLLM (kwon2023efficient) introduces PagedAttention for efficient memory management. Modern parallelization strategies (shoeybi2019megatron; aminabadi2022deepspeed) enhance the performance with multiple GPUs. However, these approaches depend on data-center GPUs connected via specialized interconnects (InfiniBand, NVLink) with throughput exceeding hundreds of GB/s (NVLink)—speeds unattainable over wide area networks (WANs).
Speculative decoding. Speculative decoding (leviathan2023fast; chen2023accelerating) reduces latency by having a smaller auxiliary model generate multiple candidate tokens for parallel verification by the main model. The process involves three phases: drafting candidates, verification, and reconciliation for generated tokens. Recent advances have focused on more efficient drafting methods, either using lighter auxiliary models or the target model with reduced parameters (bhendawade2024speculative; cai2024medusa; li2024eagle; stewart2024n; zhang2023draft).
Tree-based speculative decoding. Standard speculative decoding suffers from exponentially declining acceptance rates as sequence length increases (leviathan2023fast; chen2023accelerating). Tree-based approaches (miao2024specinfer; chen2024sequoia; svirschevski2024specexec; cai2024medusa; sun2024spectr) address this by exploring multiple paths simultaneously. Sequoia (chen2024sequoia) and SpecExec (svirschevski2024specexec) further optimize by pruning unpromising branches.
Distributed LLM serving. Split-inference approaches like Petals (borzunov2024distributed) and Helix (mei2024helix) distribute model layers and pipeline requests across multiple devices, improving throughput over memory offloading methods (ren2021zero; pudipeddi2020training), but introduce network delays that increase latency compared to data center solutions. Several works (timor2025distributed; liupearl; mcdanel2024amusd) utilize multiple GPU devices within a server node to overlap draft and verification tasks of speculative decoding for faster inference, yet with an exchange of higher cost as they require additional server devices per query.
On-device LLM inference. Frameworks like MLC LLM (mlc-llm) and Web LLM enable fully on-device inference with smaller or quantized models. While recent approaches (svirschevski2024specexec; song2024powerinfer; xue2024powerinfer) showcase the potential of leveraging user devices, they compromise output quality and latency. In contrast, our approach retains the high-quality verification stage on the server—achieving a hybrid solution that outperforms purely centralized or fully local inference.
@@ 3 Problem and Motivation
Conventional split computing. One approach to utilizing edge GPUs is split computation across edge and server resources. Split computing has been extensively studied in machine learning literature (kang2017neurosurgeon; zhou2019edge; wang2020convergence), predating transformer architectures. Traditionally, this approach divides neural network layers between server and edge, reducing server computational load or minimizing network communication by transmitting intermediate tensors rather than raw input data. However, this method is poorly suited for LLM serving due to the unique characteristics of transformers (vaswani2017attention):
Limitation #1: Excessive latency. While layer-wise splitting has been used in distributed peer-to-peer systems (borzunov2024distributed; mei2024helix) to enable serving large models otherwise infeasible on consumer GPUs, these approaches prioritize feasibility over performance. This is because, unlike earlier ML applications (he2016deep; redmon2016you; diwan2023object) that typically involve only a few communication rounds, LLM applications require communication for every generated token due to their autoregressive nature. In our distributed network setting (§5.3), layer-wise splitting increases latency by 2.35x compared to the server-only solution.
Limitation #2: Ineffectiveness in reducing I/O. LLM inference is inherently constrained by memory I/O (pope2023efficiently), as generating a single token requires accessing the entire model’s parameters. Transformer architectures, with operations like general matrix-vector multiplication (GEMV), have low arithmetic intensity relative to their memory I/O demands. While layer-wise split computing reduces computational load by offloading layers to the edge, it fails to alleviate the I/O-bound nature of inference. A more effective approach is batch-verifying multiple candidate tokens, which significantly reduces the I/O overhead per token and increases GPU throughput.
@@ 4.1 Disaggregated LLM Decoding
In our novel design, edge GPUs generate candidate tokens and send them to the server, which verifies them in a single forward pass. The server returns both the verified tokens and one additional token to the edge device, which then updates its sequence, KV cache, and continues drafting. This cycle repeats until an end-of-sequence token is generated. Critically, this approach preserves the exact output distribution of the server model (leviathan2023fast)—even when using different models for drafting and verification, the final output is guaranteed to be sampled from the same distribution as if generated by the server model alone.
@@ 4.2 Proactive Draft Generation at the Edge
Initial drafting. SpecEdge employs state-of-the-art tree-based drafting (svirschevski2024specexec). However, SpecEdge is designed to be future-proof and not tied to any specific drafting technique. It simply requires a draft phase that produces candidate tokens and is compatible with various speculative decoding approaches, including tree-based methods (chen2024sequoia; miao2024specinfer), lossy/lossless methods, and earlier single-sequence schemes (leviathan2023fast; chen2023accelerating). As the field evolves, any advances in speculative decoding can be seamlessly integrated.
@@ 5 Evaluation
Implementation and Setup. Our system’s edge-assisted configuration utilizes a server-side NVIDIA A100 GPU connected to multiple edge-side NVIDIA RTX4090 GPUs over a wide-area network. The number of RTX4090 GPUs scales with the number of concurrent requests (batch size x 2). In our experiments, we measured an average round-trip time (RTT) of 14.07ms between the local edge node and our Google Cloud instance. We conducted evaluations across various models and datasets under diverse operating conditions. The code is available at https://github.com/kaist-ina/specedge
Baseline and metrics. Our primary baseline is a server-only configuration employing tree-based speculative decoding, supplemented by autoregressive decoding and a layer-split approach that offloads part of the LLM’s layers to an edge device. SpecEdge can leverage either client-side or edge GPUs; in this evaluation, we assume it uses consumer-grade GPUs from edge cloud providers. Based on provider pricing (GCP; Vastai), the server-side A100 40GB GPU costs \$4.05 per hour, while running the 32B model requires an A100 80GB at \$5.05 per hour. In comparison, SpecEdge adds \$0.35 per hour for each RTX 4090 used. Our key metrics include cost efficiency (generated tokens per dollar), server throughput (generated tokens per unit time) and inter token latency (user-perceived output latency). All configurations produce identical output distributions as they use the same underlying models.
Models and data sets. We use four different LLMs: Qwen3-32B/14B (qwen3), Vicuna-33B (chiang2023vicuna) and Llama2-13B-chat-hf (touvron2023llama). Unless specifically noted, all models are configured with a temperature setting of 0.7. For the draft models, we use five different models: Qwen3-1.7B/0.6B (qwen3), Sheared Llama-1.3B (xia2023sheared), Tiny Llama-1.1B (zhang2024tinyllama), and JackFram-160M (miao2024specinfer). Finally, we use SpecBench (xia-etal-2024-unlocking), C4 (en) (raffel2020exploring), OpenAssistant conversations datasets (kopf2024openassistant), WikiText-2 (merity2016pointer), and MTBench (zheng2023judging).
@@
[…]
@@ 5.1 End-to-end Performance and Cost-efficiency
Beyond throughput improvements, Figure 6 demonstrates that SpecEdge also reduces inter token latency by an average of 11.24% compared to server-only speculative decoding across various batch sizes. This dual benefit—higher throughput with lower latency—is particularly valuable for interactive LLM applications where both resource efficiency and user experience are critical.
We extend our evaluation to other models, using Vicuna-33B and Llama2-13B-Chat as target LLMs across diverse datasets including C4, OAsst, WikiText-2, and MTBench. The results show that proactive edge drafting SpecEdge consistently improves performance, with greater benefits observed when the drafting models generate deeper drafts with better alignment toward the target LLMs. This alignment quality directly correlates with improvements in server throughput and reduced inter token latency. The complete results for all combinations of models and sets are provided in the Appendix B.1.
@@ 5.2 Component-wise Benefit
However, simple disaggregation alone can increase latency and leave the server underutilized (§4.1). Figure 9 compares three progressive implementations with a 14B/1.7B model pair: basic disaggregation, disaggregation with proactive edge drafting (§4.2), and complete SpecEdge with pipeline-aware scheduling (§4.3). Basic disaggregation achieves only 32.76 tokens/s throughput with higher latency, while adding proactive drafting reduces inter token latency to 28 ms. The complete SpecEdge with pipeline-aware scheduling dramatically increases server throughput to 67.89 tok/s (2.07× improvement) while maintaining latency, demonstrating the complementary benefits of each component.
@@ 5.3 System Sensitivity Analysis
Network RTT sensitivity. We evaluate the average inter token latency of SpecEdge against layer-split and server-only configurations across varying network round-trip times (RTTs), using Qwen32B on SpecBench. Layer-split configuration runs autoregressive decoding, where one-quarter of model layers run on an edge RTX4090, with the remainder on a server-side A100. The server-only configuration, unaffected by network RTT, runs tree-based speculative decoding entirely on an A100. Figure 12 shows that SpecEdge provides lower inter token latency than the server-only baseline below 50 ms RTT, with a 13.90% gain at 15 ms RTT (36.47 ms vs. 42.36 ms). Even at 65 ms, SpecEdge ’s latency rises by only 22.00% over its 15 ms RTT performance (to 44.47 ms), remaining competitive. By contrast, layer-split is much slower: at 15 ms RTT, it is 2.73× slower than SpecEdge (99.63 ms vs. 36.47 ms), increasing to 3.35× slower performance at 50 ms RTT (134.90 ms vs. 40.26 ms). This resilience stems from SpecEdge ’s less frequent communication rounds and proactive edge drafting, which offset network latency more effectively than layer-split approaches.
@@
Performance with varying edge devices. Figure 12 presents the speedup achieved by SpecEdge when the server is assisted by different edge devices (RTX 4090, 4070 Ti Super, 3090, and 2080 Ti) across three target-draft model combinations. The speedup is measured relative to default autoregressive decoding using only the A100 server GPU. The consistent speedup across all model combinations confirms the architecture’s robustness to different hardware configurations. As expected, more powerful edge GPUs like the RTX 4090 deliver greater speedups, while even more affordable options like the RTX 2080 Ti still provide significant acceleration. This demonstrates that SpecEdge’s approach remains effective across a spectrum of edge hardware capabilities, allowing deployment flexibility. Additional results with lighter GPUs (3060 Ti, and 2080 Ti) are available in Appendix B.2.
@@
Performance with alternative drafting approaches. We also evaluated SpecEdge with a non-tree speculative decoding approach (leviathan2023fast) to demonstrate versatility beyond tree-structured methods. Using this alternative architecture, SpecEdge achieved up to 1.96× higher server throughput (Llama2 13B/TL 1.1B pairing) and 1.67× better cost efficiency compared to server-only deployments. Performance gains remained consistent across different model combinations, with even the smallest draft model (JF 160M) delivering a 1.52× throughput improvement while preserving end-user speedup. Complete results are available in Appendix B.3.
Performance with batch drafting method. We explore an alternative drafting configuration where a single edge GPU serves concurrent requests through batching. This approach enables operators to reduce the number of edge GPUs. Experiment with various batch sizes revealed a trade-off between better cost efficiency and increased latency. This configuration could be advantageous in budget-constrained deployments where latency tolerance is higher. Full results across model combinations are available in Appendix B.4.
Performance under reasoning mode. Modern LLMs provide reasoning mode for enhancing output quality, where reasoning tokens might influence speculative decoding efficiency (wei2023chainofthought; qwen3). To explore SpecEdge performance under reasoning mode, we measured accepted tokens, server throughput, and cost efficiency with and without reasoning enabled. Results in Appendix B.5 show consistent improvements across all three metrics when reasoning is active, implying that our system inherently benefits from the redundancy in reasoning processes, which enhances speculative decoding performance.
Cost analysis with various GPU Providers. To ensure SpecEdge’s cost-efficiency findings generalize beyond a single provider, we validated SpecEdge’s performance across diverse cloud environments. We compared results using GPUs from multiple GPU providers (Vast.ai (Vastai), Runpod (runpod), and TensorDock (TensorDock)) and Cloud Service Providers (Google Cloud Platform (GCP), Amazon Web Services (AWS), and Microsoft Azure (Azure)), accounting for the pricing variations. Across all tested configurations, SpecEdge consistently delivered cost efficiency improvements, confirming that our architectural benefits persist regardless of the specific cloud infrastructure. Detailed cross-provider comparisons are presented in Appendix C.
@@
Detailed Case Study. Complementing our quantitative evaluation, Appendix D offers a visualization of SpecEdge in action. Using Llama models responding to a query about Dyson Spheres, we trace the complete token generation lifecycle—from initial drafting through verification and subsequent accelerated generation. The case study specifically highlights two key operational advantages: (1) how edge GPUs remain productive during server verification phases through proactive expansion, and (2) how successful draft alignments lead to deeper draft candidates in subsequent rounds.